Overhead Variances, Four-Variance Analysis

Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 122,000 units requiring 488,000 direct labor hours. (Practical capacity is 508,000 hours.) Annual budgeted overhead costs total $795,440, of which $566,080 is fixed overhead. A total of 119,100 units using 486,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $261,000, and actual fixed overhead costs were $555,150.

Required:

1. Compute the fixed overhead spending and volume variances.

Fixed Overhead Spending Variance $
Fixed Overhead Volume Variance $

2. Compute the variable overhead spending and efficiency variances. Do not round intermediate calculations

Variable Overhead Spending Variance $
Variable Overhead Efficiency Variance $
Answer #1

Ans: Fixed overhead spending variance= Actual fixed overhead- budgeted fixed overhead

= $555,150- $566080

=$ 10930(favorable)

Fixed overhead volume variance = Standard cost for actual ouptut- budgeted cost

= 4.64**x119100)-566080

=552624-566080

= -$13456(unfavorable)

** Standard cost= 566080/122000=4.64

2). Variable overhead spending variance=actual variable cost- standard cost for actual ouptut

= 261000-(0.47*488000)

=261000-229360

=$31640(unfavorable)

** Standard variable cost= $795440-566080/488000=0.47 per hour

B). Variable overhead efficiency variance= actual hours* standard per hour- standard hours* standard rate per hour

= 480000*0.47-(476400*0.47)

=225600-223908

=1692(unfavorable)

* Standard rate per hour= Budgeted variable overhead/ Standard hours

=( 795440-566080)/488000

=0.47 per hour

* Standard hour for actual output if 488000 hours are required to produce to make 122000 units then to make 119100 units:

488000*119100/122000=476400 hours required